Buoyancy is an interesting phenomenon which defines about an object which has tendency to float in water. The density difference between the water and atmospheric air directly affects the buoyance effects. In a typical experimental study, an object weighs 392.4 N in atmospheric air and 196.2 N in water.

Determine the volume of body and its specific gravity.



1
Expert's answer
2021-07-06T11:51:55-0400

upthrust force=weight in air-weight in water=392.4N-196.2N=196.2N

"upthrust=" weight of water displaced="density\times volume\times gravity"

"196.2=1000\times volume\times 9.8"

"volume =\frac{196.2}{1000\times 9.8}=0.02m^3"

volume of body"=0.02m^3"

mass of the body"=w\/g=392.4\/9.8=40.0408kg"

density of the body"=mass\/volume=40.0408kg\/0.02m^3=2002.04kgm^{-3}"

specific gravity"=" density of body"\/" density of water"=2002.04\/1000=2.0020"